OpenStudy (anonymous):

What is the length of the midsegment AB? Trapozoid MNOP. M is at 8 comma 6, N is at 17 comma 6, O is at 20 comma 0, and P is at 0 comma 0; midsegment AB is located such that A is on side MP and B is on side NO A. 22 B. 31 C. 14.5 D. 29

OpenStudy (jdoe0001):

hmmm how long is MN? how long is PO?

OpenStudy (anonymous):

That is what i do not know @jdoe0001

OpenStudy (jdoe0001):

well.. look at the picture, look at the coordinates, what would be the difference between M and N? and the difference between P and O ?

OpenStudy (anonymous):

Well MN is shorter and PO is longer. It only gives me points not numbers so i do not know the length of them.

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

those are numbers on the "x" and "y" axis.... so anyhow from (8, 6) to (17, 6) will be well 17-8

OpenStudy (jdoe0001):

and P is at (0, 0) and O is at (20, 0) so 20 -0

OpenStudy (jdoe0001):

anyhow that'd give you how long those 2 segments are, and those parallel lines to each other in a trapezoid are called the BASES, and then \(\bf \textit{MidSegment of a Trapezoid}= \cfrac{1}{2}\times (base1+base2)\)

OpenStudy (anonymous):

I am still not getting it :/ Sorry @jdoe0001 So base 1= 20-0 and base2 = 17-8. But how do you add them together?

OpenStudy (jdoe0001):

hmmm well. you need to get the length of the 2 parallel lines in the trapezoid

OpenStudy (jdoe0001):

the 2 parallel lines in the trapezoid are PO and MN, thus those are the bases

OpenStudy (anonymous):

I am getting really frustrated. I really don't get this. My tutor doesn't do anything either. :/ @jdoe0001

OpenStudy (anonymous):

@jdoe0001 are you there?